### v2

# Desynchronization Index: A New Connectivity Approach for Exploring Epileptogenic Networks

Abstract
Stereoelectroencephalography (SEEG) is an invasive surgical procedure to record the electrical activities in cortical brain regions, aiming at identifying the Epileptogenic Zone (EZ) in patients with drug-resistant epilepsy. To improve the accuracy of the EZ definition, SEEG analysis can be supported by computational tools, among which the Epileptogenic Index (EI) represents the most common solution. However, the scientific community has still not found an agreement on which quantitative biomarkers fully characterize the cortical sites within the EZ. In this work, we design a new algorithm, named Desynchronization Index (DI), to assist neurophysiologists in SEEG interpretation. Our framework hypothesizes that the EZ is identified by the cortical sites that get abnormally desynchronized from the network during the seizure onset. We test the proposed method over a SEEG dataset of 11 patients, comparing its accuracy in terms of EZ localization against the EI algorithm and clinical ground truth. Our results indicate that the DI algorithm underscores specific connectivity dynamics that can hardly be identified with a pure visual analysis, increasing system sensitivity in EZ localization with respect traditional methods.
Index Terms:
Stereoelectroencephalography, drug-resistant epilepsy, cortical connectivity, time-varying networks, phase transfer entropy.
This paper has been submitted to IEEE Transactions on Neural Systems and Rehabilitation Engineering
I Introduction
The localization of the Epileptogenic Zone (EZ) is the fundamental prerequisite for performing epilepsy surgery in people with drug-resistant epilepsy [1]. Ensuring a higher accuracy in the definition of the EZ makes it possible to reduce the extent of the surgery, mitigating the risk of complications or chronic deficits. In the most complex scenario, the EZ can be identified only through invasive procedures, among which Stereoelectroencephalography (SEEG) represents the most accurate methodology [2]. An SEEG recording involves the surgical implantation of electrodes into the patient’s brain, enabling the monitoring of electrical activity from deep cortical regions, which is essential for the identification of the EZ. As a result, SEEG-guided surgical resections lead more than $60\%$ of patients to be seizure-free after the intervention [3]. However, SEEG has the drawback of providing very focused information: if none of the intracranial electrodes intersects the EZ, the procedure may result in an unclear EZ definition and an unsuccessful epilepsy surgery [4].
Nowadays, SEEG interpretation is mainly based on visual analysis, focusing on the early phase of the seizure discharge. The final goal is to define an EZ that can be focal or distributed among multiple cortical structures, a scenario falling under the concept of epileptogenic networks [5]. The above process is particularly complex since SEEG signals include hundreds of components, whose analysis requires the involvement of highly specialized neurophysiologists. For this reason, in the last years, SEEG interpretation has started being supported by computational tools apt at characterizing the EZ and, more in general, the dynamics leading to seizure generation [6]. The most common methods analyze the SEEG power spectrum, searching for the cortical sites generating fast oscillations, i.e., electrical activities with frequencies in the gamma range ($[30,100]$ Hz), as done in the cases of Epileptogenic Index (EI) [7] and Epileptogenic Maps (EMs) [8].
Other approaches try to model the cortical connectivity, identifying which brain structures show abnormal connectivity behavior during the seizure onset. In this context, the first studies have focused on the electric responses generated through intracranial stimulations, known as Cortico-Cortical Evoked Potentials (CCEPs) [9]. The SEEG connectivity can also be estimated by analyzing the phase distribution of the electrical signals generated by each cortical site, as performed by the Phase Locking Value (PLV) and the Phase Lag Index (PLI), avoiding biases associated with amplitude differences.
Although multiple quantitative approaches have been developed for analyzing the connectivity dynamics of epileptogenic networks, the literature often presents discordant findings. Most works agree that the ictal phase (i.e., the period involving the seizure discharge) is preceded by a connectivity reduction, followed by the opposite phenomenon during the seizure propagation [10]. Recently, it has been hypothesized that the EZ presents an abnormal inward information flow during the inter-ictal phase [11], and the epileptogenic level of each cortical site can be assessed by the connection exerted during the epileptic discharge [12].
In this work, we present a new framework for supporting neurophysiologists in SEEG interpretation. Specifically, we develop a connectivity-based algorithm, named Desynchronization Index (DI), that exploits the Phase Transfer Entropy (PTE) to estimate the relations between brain structures and detect connectivity anomalies in the resulting time-varying network. According to our approach, the EZ is identified by the cortical sites that present an independent behavior during the ictal phase. Overall, the main contributions of this manuscript consist of the following points:
  • we design a new connectivity-based metric, named desynchronization level, describing the tendency for a cortical site to be disconnected from the rest of the brain;
  • we design a new algorithm, named DI, which estimates the epileptogenic degree of cortical sites according to their connectivity dynamics, quantified in terms of desynchronization level;
  • we implement the DI algorithm to analyze seizure data in $11$ consecutive patients with drug-resistant epilepsy who underwent SEEG monitoring;
  • we analyze the clinical utility of our method and its main disadvantages and drawbacks in comparison to the EI algorithm and clinical ground truth.
Our results indicate that the combined use of the DI and EI algorithms can offer important support to neurophysiologists for the identification of the EZ. In particular, the DI algorithm underscores anomalous connectivity dynamics which can hardly be detected by adopting a pure visual analysis or by using EI as a standalone tool. Thus, the proposed technique may be used to identify cortical regions that, despite not presenting fast oscillations, surround the EZ and play an active role in the generation and propagation of seizures.
II Related Work
In clinical practice, the interpretation of SEEG recordings is mainly based on visual analysis, which is a highly time-consuming process because of the high number of signals that must be reviewed [13]. The most common marker of the seizure onset is given by low-voltage fast activities, i.e., electrical discharges spanning beta frequencies ($[13,30]$ Hz), usually observed in mesial temporal seizures, and gamma frequencies ($[30,100]$ Hz), observed in neocortical seizures. The quantification of fast activities and the network dynamics leading to seizures can be achieved by analyzing the SEEG spectrum. Such an approach is followed by the EI algorithm [7], which represents the most routinely used method to assist clinicians in SEEG interpretation [14].
The EI algorithm, occasionally combined with other biomarkers [15], has been recognized to provide the best accuracy in terms of EZ localization and surgical prediction. In this regards, the epileptogenicity rank extends the EI framework by assigning weights to SEEG sites depending on to their distance from the hypothesized center of the EZ [16]. This technique has the obvious limitation of biasing the outcome of the quantitative analysis by a prior assumption. Hence, the epileptogenicity rank reduces the number of false positives when the clinical hypothesis is correct, but dramatically decrease the algorithm sensitivity in the other case.
While EI focus on the energy distribution in the frequency domain, other approaches try to estimate the variation in neuronal connectivity during the ictal phase. Several tools have been proposed for such a goal, including the PLV [17] and the PLI [18], which involve different trade-offs in terms of sensitivity and precision [19]. In particular, the PLI has the advantage of avoiding the noisy information associated with the volume conduction [20], reducing the number of false positives. A more advanced method is the PTE [21], which exploits the Granger Casuality (GC) [22] to estimate both the delay and the magnitude of neuronal connections, representing a very promising technique for SEEG analysis [23].
Alongside designing the EI, Bartolomei and colleagues analyze SEEG data by the nonlinear regression coefficient (h2) [24], and show that the epileptogenic network is associated with reduced connectivity in the period that foresees the ictal phase  [25]. While the h2 coefficient is agnostic to the signal frequencies, most connectivity models discern cortical relationships depending on the frequency bands. Following this principle, the authors of [26] estimate connectivity via linear correlation and observe that seizure frequency is proportional to the EZ outward connectivity. Similar methodologies are given in [27], whose authors combine pairwise correlation with Euclidean distance to estimate the EZ extension, and [28], where graph theory is exploited to identify the network hubs most contributing to the seizure propagation.
Recently, the use of Spearman rank correlation [29] to assess the coupling between brain structures has shown good results for both the detection of the EZ and the prediction of surgical outcomes [30]. Other connectivity-based approaches correlate the phase of slow oscillations ($<30$ Hz) with the amplitude of fast oscillations ($>30$ Hz), as occurred for the Phase Slope Index (PSI) and the Phase Amplitude Coupling (PAC) [31]. Exploiting the PSI, [32] shows that the EZ presents higher outward connectivity during the seizure onset than the resting state. Similar results are observed in [33], where the PAC between high (gamma) and low (delta and theta) frequency bands is computed. The authors of [34] seem to contradict these findings and, using the Partial Direct Coherence (PDC) [35] as a connectivity measure, assess that the EZ presents higher inward connectivity during seizure generation.
In the last few years, the popularity of Machine Learning (ML) has encouraged many researchers to apply data-driven techniques to epileptogenic networks. For instance, the authors of [36] exploit hypergraph learning for automatically detect SEEG fast oscillations, while [37] investigates different patient-specific ML models for analyzing scalp electroencephalography signals. Besides, [38] exploit a Support Vector Machine (SVM) model to integrate multiple electrophysiological features, including both fast oscillations and low frequency suppression, for enhancing EZ detection. This supports the fact that ictal channels are not identified by specific frequency components, and more holistic methodologies are required to describe seizure generation. On the other hand, [38] and similar works consider as ground truth the area resected by epilepsy surgery, which is usually far larger than the EZ to maximize the probability of seizure reduction. In general, the high complexity of SEEG analysis makes it is extremely hard to build supervised datasets, which prevents ML solutions from obtaining reliable performance in this field [39].
Although the large variety of works implementing connectivity-based metrics for SEEG interpretation, results from the literature do not lead to an agreement on which biomarkers characterize the EZ. Initial works show that the ictal transition is characterized by a connectivity reduction, but the EZ is often identified by the cortical sites with a higher coupling during the seizure propagation. In this work, we try to overcome such questions, by quantifying the epileptogenicity level of SEEG sites in terms of desynchronization. With this latter, we denote the tendency of a cortical site to assume an independent electrical behavior that cannot be inhibited by neighboring brain structures.
III Analytical Model
In this section, we first describe how SEEG signals are modeled within our framework and introduce the EI algorithm, which represents the state-of-the-art for detecting the EZ. Then, we present the connectivity model used to infer the relation between SEEG sites starting from their phase distribution. Finally, we introduce the DI algorithm, which evaluates the epileptogenicity level of each site according to its tendency to desynchronize from the network.
III-A Epileptogenic Index
We model a single SEEG recording as a multidimensional signal with $N$ different components, named channels, one for each cortical site analyzed. In the following, we denote by $\mathcal{N}$ the set of channels and by ||$\cdot$|| ⋅ | the cardinality operator, so that the number of cortical sites is equal to $N=|\mathcal{N}|$. According to our approach, each channel $x\in\mathcal{N}$ is segmented in multiple overlapping windows $x(t)$, with $t=0,\Delta t,2\Delta t,…$, where $\Delta t$ represents the time-shift between consecutive windows. Given the window duration $T_{\text{window}}$ and the sampling frequency $f_{s}$, the sample number per window is $n=T_{\text{window}}/f_{s}$.
The EI algorithm requires to evaluate the signal energy in the frequency domain [7]. Given a specific window $x(t)$, the algorithm first computes the Fourier Transform (FT) of $x(t)$, obtaining a complex value $X(t,f)$ for each frequency in $[0,f_{s}/2]$. Hence, the energy ratio between the high- and low-frequency bands of the target signal $x$ is computed as
[…]
To avoid the underlined issue, the EI algorithm considers a CUmulative SUM (CUSUM) control chart to discern the channels with abrupt increases in the energy ratio. Given a sequence of observations $\omega(t)$, interspersed by a period $\Delta t$, the CUSUM control chart is defined by the function
$\Gamma(\omega,t)=\begin{cases}\max\left\{0,\Gamma(\omega,t-\Delta t)+\frac{\omega(t)-\mu_{\omega}}{\mu_{\omega}}-\gamma\right\},&t>0;\\ \omega(t),&t=0;\end{cases}$ (2)
where $\gamma$ is a tuning parameter that makes the statistic less or more sensitive to the new observations, while $\mu_{\omega}$ and $\sigma_{\omega}$ are the estimates of the mean and standard deviation of $\omega$. In the case of the EI algorithm, the observations $\omega(t)$ are given by the energy ratio $E_{x}$ associated with each channel $x\in\mathcal{N}$.
The original version of the EI algorithm does not consider the standard deviation $\sigma_{E_{x}}$ in the normalization and dynamically re-estimates the mean $\mu_{E_{x}}$ every time the energy ratio varies significantly. Instead, in this work, we propose to estimate both the $\mu_{E_{x}}$ and $\sigma_{E_{x}}$ statistic by looking at the period immediately preceding the ictal discharge. Practically, for each channel $x\in\mathcal{N}$, we compute $\mu_{E_{x}}$ and $\sigma_{E_{x}}$ over the time interval $[t_{\text{base}},t_{\text{start}}]$, where $t_{\text{start}}$ is the instant at which the epileptic discharge starts to form. Then, we compute the cumulative sum $\Gamma(E_{x},t)$ of the energy ratio $E_{x}$ for each time $t\in[t_{\text{start}},t_{\text{end}}]$, where $t_{\text{end}}$ is the instant at which the epileptic discharge propagates within the overall network.
Given a channel $x\in\mathcal{N}$ and $\Gamma(E_{x},t)$, $\forall$ $t\in[t_{\text{start}},t_{\text{end}}]$, we define the activation time $t_{E_{x}}$ of the channel as the time instant at which $\Gamma(E_{x},t)$ reaches the highest value:
$t_{E_{x}}=\operatorname*{arg\,max}_{t\in[t_{\text{start}},t_{\text{end}}]}\Gamma(E_{x},t)$ (3)
Besides, we define the tonicity $c_{E_{x}}$ of the channel as the sum of the energy ratio $E_{x}$ computed over the interval following the channel’s activation time:
[…]
where $\delta$ is a tuning parameter determining the interval over which $c_{E_{x}}$ is computed.
Hence, we can select the $M$ channels with strongest energy ratio changes by defining the set $\mathcal{N}_{E}\subset\mathcal{N}$, of cardinality $|\mathcal{N}_{E}|=M$, whose elements $y$ comply with the conditions
$\Gamma(E_{y},t_{E_{y}})>\Gamma(E_{x},t_{E_{x}}),\;\forall\,x\in\mathcal{N}\,\backslash\,\mathcal{N}_{E}.$ (5)
Under this perspective, $M$ denotes the maximum number of epileptogenic channels that are detected by the algorithm. The EI value of each channel $x\in\mathcal{N}_{E}$ is then given by
$EI_{x}=\frac{c_{E_{x}}}{t_{E_{x}}-t_{\text{start}}},$ (6)
while the EI value of each channel $x\in\mathcal{N}\,\backslash\,\mathcal{N}_{E}$ is set to 00. We observe that $EI_{x}$ is proportional to $c_{E_{x}}$ and decreases as a function of the time difference between the channel activation time $t_{E_{x}}$ and the beginning of the ictal phase $t_{\text{start}}$.
The above formulation stands out from the original EI algorithm by reducing the number of parameters to be manually set. In particular, $t_{\text{start}}$ and $t_{\text{end}}$ derive from the clinical interpretation of the SEEG recording and denote the period during which the epileptic discharge takes shapes. Instead, $t_{\text{base}}$ should be defined in such a way to ensure that the statistics $\mu_{E_{x}}$ and $\sigma_{E_{x}}$ are as accurate as possible. Hence, the only parameters to be tuned are the maximum number $M$ of epileptogenic channels, the interval $\delta$ over which the tonicity is computed (used in (4)), and the weight $\gamma$ of the CUSUM control chart (used in (2)), which trades off between the false alarm and miss-detection probabilities.
III-B Effective Connectivity Model
To estimate the SEEG connectivity dynamics, we consider the PTE algorithm, which, given a couple of channels $x,y\in\mathcal{N}$ at time $t$, models their coupling according to the phase distributions of $x(t)$ and $y(t)$. From this perspective, PTE may lead to different results depending on how we model the phase distribution. Given a specific window $x(t)$, we first compute its analytic representation $\mathcal{X}(t)=x(t)+HT\left(x(t)\right)$, where $HT(\cdot)$ is the Hilbert transform. Notably, $\mathcal{X}(t)$ is associated with $n=T_{\text{window}}/f_{s}$ phase values, representing the evolution of the instantaneous phase $\theta_{x}(t)$ of $x(t)$ in time. According with the Sturges rule [40], we can infer the bin width $\vartheta$ of $\theta_{x}(t)$ as $\vartheta=2\pi/(\log_{2}(n)+1)$.
The proposed methodology quantizes the phase values to reduce the influence of volume conduction on the connectivity model. A similar same principle is adopted by the PLI algorithm, which, however, does not consider the effectiveness of the network connection [18]. In addition, PLI presents critical pitfalls, including underestimating those connections associated with small propagation delays or with signals whose phase difference distribution fluctuates around zero. Our technique tries to act as a compromise between PLI and PLV, which, notably, suffers the effect of volume conduction and leads to many false connectivity phenomena.
Lets denote by $\theta_{x}(t)$ and $\theta_{y}(t)$ the phase distribution of $x(t)$ and $y(t)$, respectively. According to the PTE algorithm, the connectivity that the channel $x$ exerts on the channel $y$ at time $t$, considering a lag $\tau$, is given by
$\begin{gathered}T_{x\rightarrow y}(t,\tau)=H\left(\theta_{y}(t),\theta_{y}(t+\tau)\right)+\\ H\left(\theta_{x}(t),\theta_{y}(t)\right)-H\left(\theta_{x}(t),\theta_{y}(t),\theta_{y}(t+\tau)\right)-H\left(\theta_{y}(t)\right),\end{gathered}$ (7)
where $H(\cdot)$ denotes the entropy function. The PTE formulation, as reported in (7), is a directed connectivity measure, which means that, in general, $T_{x\rightarrow y}(t,\tau)\neq T_{y\rightarrow x}(t,\tau)$. In order to remove the dependency from $\tau$, with a slight abuse of notation, we redefine the phase transfer entropy between $x(t)$ and $y(t)$ as the maximum value of $T_{x\rightarrow y}(t,\tau)$ among multiple lags in the set $[0,\tau_{max}]$:
$T_{x\rightarrow y}(t)=\max_{\tau\in[0,\tau_{max}]}T_{x\rightarrow y}(t,\tau).$ (8)
By doing so, we obtain that the magnitude of the effective connection exerted on the channel $y$ by the channel $x$ at time $t$ is given by $T_{x\rightarrow y}(t)$, while the propagation delay associated with such a connection is:
$\tau_{x\rightarrow y}(t)=\operatorname*{arg\,max}_{\tau\in[0,\tau_{max}]}T_{x\rightarrow y}(t,\tau).$ (9)
Hence, if $\tau_{x\rightarrow y}(t)=0$, there is a zero-propagation delay for the information flow going from $x(t)$ to $y(t)$.
III-C Desynchronization Index
To identify ictal channels from the connectivity analysis, we follow a similar approach to that presented in [41], where Sparks and colleagues extended the Exponentially Weighted Moving Average (EWMA) statistic to detect anomalous behaviors in time-varying networks. In particular, we assume that the generation of epileptic discharges is associated with an abrupt reduction in the communication between the EZ and the rest of the cortical structures.
Given a sequence of observations $\omega(t)$, interspersed by a period $\Delta t$, the EWMA control chart is defined by the function:
$\mathcal{E}(\omega,t)=\begin{cases}\alpha\omega(t)+(1-\alpha)\mathcal{E}(\omega,t-\Delta t)\;&t>0;\\ \omega(t)\;&t=0;\end{cases}$ (10)
where $\alpha\in[0,1]$ is a tuning parameter that determines the rate at which older information influences the EWMA statistic. In particular, as $\alpha\rightarrow 1$, the function gives more importance to the most recent observations, better capturing rapid process changes at the cost of an increased false alarm probability.
In the following, we write $\mathcal{T}(t)$ to indicate the distribution of the inter-channel connections $T_{x\rightarrow y}(t)$, $\forall\,x,y\in\mathcal{N}$, at time $t$. Hence, we denote by $E[\mathcal{T}(t)]$, $P_{75}[\mathcal{T}(t)]$, and $P_{25}[\mathcal{T}(t)]$ the expectation, the $75$th percentile, and the $25$th percentile of $\mathcal{T}(t)$. We use the EWMA algorithm, as defined in Eq.(10), to compute a smoothed version of the above statistics as well as of each connection $T_{x\rightarrow y}(t)$, $\forall\,x,y\in\mathcal{N}$. Then, we define $\mathcal{N}_{x}^{\text{in}}(t)$ as the set of channels presenting abnormal inward connectivity with respect $x$ at time $t$:
[…]
In other words, we consider abnormal those connections that go outside the inter-quartile range of the time-varying network distribution. This, on the one hand, allows us to implicitly customize the framework to each patient without the need to define additional parameter as in [41]. On the other hand, choosing other percentile ranges for discerning abnormal connections could lead to different trade-offs in terms of accuracy, and, in the future, further investigation of how to tune such a parameter is required.
TABLE I: Demographic and clinical data.
Patient index Sex Age (SEEG) Age (onset) Seizure frequency EZ localization PZ localization Surgery Outcome (1 year) Histology
1 Male 39 33 Weekly Right temporal Temporal pole Lobectomy Seizure free HS 1
2 Male 24 1 Daily Left orbito-frontal Fronto-insular RF-TC 1 seizure in 1 year Not available
3 Male 36 16 Monthly Left temporal Temporal pole Lobectomy Seizure free FCD 1b
4 Female 52 3 Weekly Right temporal Temporal and FOI Lobectomy Seizure free HS 1
5 Female 25 2 Weekly Right temporo-basal Temporo-parietal and insular Lobectomy Seizure free Aspecific
6 Male 48 22 Weekly Right temporal Orbito-frontal, and insular Lobectomy Seizure free HS 1
7 Male 36 9 Weekly Right temporo-occipital Temporo-basal and mesial RF-TC Not available Not available
8 Female 49 12 Monthly Left temporal Temporo-lateral Lobectomy Not available HS 1
9 Female 38 13 Monthly Right temporo-occipital Temporo-mesial RF-TC Not available Not available
10 Male 32 1 Monthly Left temporo-occipital Parietal and insulo-opercular RF-TC Not available Not available
11 Male 24 18 Weekly Right temporal Temporal pole RF-TC Seizure free Tumor
We define the actual $\psi_{x}^{\text{in}}(t)$ and expected $\hat{\psi}_{x}^{\text{in}}(t)$ inward connection density of channel $x$ at time $t$ as follow:
$\psi_{x}^{\text{in}}(t)=\sum_{y\in\mathcal{N}_{x}^{\text{in}}(t)}\mathcal{E}(T_{y\rightarrow x},t);$ (12)
$\hat{\psi}_{x}^{\text{in}}(t)=|\mathcal{N}_{x}^{\text{in}}(t)|\cdot\mathcal{E}(E[\mathcal{T}],t).$ (13)
We can now define the desynchronization level of a channel $x\in\mathcal{N}$ at time $t$ as
$D_{x}(t)=\sqrt{\hat{\psi}_{x}^{\text{in}}(t)}-\sqrt{\psi_{x}^{\text{in}}(t)}.$ (14)
In particular, at each time $t$, $D_{x}(t)$ represents the tendency for channel $x\in\mathcal{N}$ to assume an independent behavior.
We apply the CUSUM chart to the desynchronization level, thus obtaining $\Gamma(D_{x},t)$ for each channel $x\in\mathcal{N}$ and for each time $t\in[t_{\text{start}},t_{\text{end}}]$. Hence, using equations (3) and (4), we can compute the activation times $t_{D_{x}}$ and the tonicities $c_{D_{x}}$, $\forall$ $x\in\mathcal{N}$, considering the desynchronization level instead of the energy ratio. Following the same approach presented in III-A, we define the set $\mathcal{N}_{D}$, including the $M$ channels showing the strongest changes in terms of connectivity. Finally, for each channel in $\mathcal{N}_{D}$, we define the DI values as
$DI_{x}=\begin{cases}\frac{c_{D_{x}}}{t_{D_{x}}-t_{\text{start}}}&x\in\mathcal{N}_{D},\\ 0&x\notin\mathcal{N}_{D}.\end{cases}$ (15)
Appreciably, the DI algorithm follows the same settings of EI, making it necessary to tune only three parameters, namely the maximum number $M$ of epileptogenic channels, the interval $\delta$, and the weight $\gamma$.
IV Evaluation Methodology
In this section, we present the clinical dataset where our tool is evaluated, describing the SEEG recording process and the patient characteristics. Then, we showcase how the EI and DI algorithms are implemented and describe how we evaluate their performance in the detection of the EZ.
IV-A Clinical Dataset
This study considers a cohort of $12$ consecutive patients that were monitored through SEEG at IRCCS Institute of Neurological Sciences of Bologna from January 2022 to June 2024. The protocol for the study was approved by the local ethics committee (protocol number 741-2021, committee code 97338), and written informed consent was obtained from each patient. In all the cases, the SEEG implant included multiple electrodes, each presenting 555518181818 recording sites, named contacts; the number and location of the electrodes were patient-tailored, while each contact was $22$ mm in length, and separated by $1.5$ m from neighboring contacts.
The SEEG implantation followed the workflow developed at Niguarda Hospital [42], which involves the construction of a multimodal scene of the patient’s brain. The scene allows for a comprehensive evaluation of all the anatomical information regarding the cortical area explored by each contact. The SEEG signals were recorded using the Nihon Kohden 2100 polygraph, considering 192 or 256 channels and a sampling frequency of $f_{s}=1000$ Hz. High-definition synchronized videos were recorded for the whole duration of each SEEG monitoring (up to 20 days per patient), enabling a correlation between electrical and clinical features. Finally, we collected the main demographic and clinical data of each patient, including age at disease onset, seizure semiology, neuropsychological examinations, neuroimaging findings, scalp EEG recordings, surgical outcomes, follow-up information, and histopathological diagnoses.
From the initial dataset, a single patient was excluded since the SEEG recording did not allowed the localization of the EZ and no epilepsy surgery was conducted. The final cohort comprises 7 males and 4 females with an average age of 36.6 years (range 24-49) at the recording time and 8.8 years (range 1-22) at the time of the disease onset. The clinical data pointed to a probable EZ in the temporal lobe in 7 cases (5 right, 2 left), in the left-frontal lobe in a single case, and in temporo-occipital lobe in the remaining cases (2 right, 1 left). At the time of SEEG implantation, seizure frequency ranged from daily to monthly, depending on the patient. Radio-frequency thermocoagulations enabled the control of seizures in a single patient, while the remaining underwent anterior-temporal lobectomy. Histopathological analysis led to clinical results compatible with hippocampal sclerosis in 4, and cortical dysplasia in 1 cases. Demographic and clinical details of the studied population are reported in Tab. I.
HS stays for hippocampal sclerosis FCD for focal cortical dysplasia, RF-TC for radio-frequency thermocoagulation, FOI for fronto-orbito-insular.
IV-B Performance Evaluation
For each patient, we consider the channels belonging to the EZ as the ground truth. The selection of the EZ channels was performed by two board-certified neurophysiologists (L.D-V. and E.P.), who were blinded to the output of our computational framework. A single seizure per patient was selected, and an SEEG epoch of $T_{\text{epoch}}=200$ seconds is analyzed for each seizure, considering the period before the complete diffusion of the epileptic discharge. If multiple seizures were available, we chose to discard the seizures recorded during the first two days of the monitoring period (during which the pharmacological treatment is still active) and those occurring during a seizure cluster.
Before running the EI and DI algorithms, the SEEG signals are pre-processed through a notch filter centered at $f_{\text{notch}}=50$ Hz. In the analysis, we exclude the sites exploring the White Matter (WM) since they are not sources of electric signals. The selection of the WM channels was performed by a board-certified neurophysiologist (L.F.), according to the electrical activity of the cortical sites and their localization in the SEEG multimodal scene. No other cortical sites are excluded from the analysis, making our framework agnostic to the specific patient and seizure to be analyzed.
The output of the EI and DI values are normalized in the $[0,1]$ interval, and a threshold $\eta$ (in the same range) is set for classifying a channel as epileptogenic. Hence, the following performance metrics are considered:
  • sensitivity (or true positive rate), which is the ratio between the number of channels correctly classified as epileptogenic by the algorithm and the total number of epileptogenic channels in the SEEG implant;
  • precision (or positive predictive value), which is the ratio between the number of channels correctly classified as epileptogenic by the algorithm and the total number of channels classified as epileptogenic by the algorithm;
  • accuracy, which is the ratio between the number of channels correctly classified (as epileptogenic or not epileptogenic) by the algorithm and the total number of channels in the SEEG implant.
The EI and DI algorithms are both implemented as standalone tools or combined according to the “and” and “or” Boolean functions. Practically, in the first case (named “EI and DI”), a channel is classified as epileptogenic only if both the algorithms mark it as positive. Instead, in the second case (named “EI or DI”), a channel is classified as epileptogenic if at least one of the algorithms classifies it as positive.
We first consider as ground truth the EZ as defined by the neurophysiologists; then, the analysis is repeated by considering a subset of the EZ channels, named Detectable Epileptogenic Zone (D-EZ). This latter includes all the EZ channels classified as positive by at least one of the algorithms (i.e., using the EI or DI approach). This approach makes it not possible to describe the algorithms’ performance in absolute terms but only to compare their results.
As the last step, we evaluate the performance of the algorithm in identifying the channels involved in the propagation of epileptic discharges. To this goal, we define a new detection class, named Detectable Propagation Zone (D-PZ), that includes all the channels classified as positive by at least one of the algorithms (i.e., using the “EI or DI” approach) and, afterward, labeled as propagator by the neurophysiologists. Although the EZ removal is the primary goal of epilepsy surgery, also the identification of propagator channels is extremely valuable for surgical planning. In the case that some epileptogenic sites cannot be removed (e.g., because this will endanger some cognitive functions of the patient), including the Propagation Zone (PZ) in the resection area may be the only solution for ensuring a seizure reduction.
We observe that other works in this field consider as ground truth the area resected after epilepsy surgery or the target of thermocoagulation treatment. However, using these information may not enables the accurate definition of the EZ as, notably, epilepsy surgery is designed to remove a larger area than the sole EZ, to maximize the probability of seizure reduction in the patient. According to our framework, instead, the definition of the EZ is based on the clinical evaluation of the same SEEG data taken as input by our computational framework. This allows us to evaluate the EI and DI algorithms in fairly context, i.e., during the clinical stage in which it is expected to be used, against a pure human analysis.
For all the above reasons, relying on neurophysiological evaluation for the definition of the clinical ground truth represents one of the main strength of this manuscript. On the other hand, this prevents us from considering other SEEG datasets to further assess the accuracy of our algorithms, as pre-existent data are likely to be associated with labels different from those used in this work. In the future, repeating the analysis involving multiple centers, following a consistent methodology for SEEG analysis, will be paramount for the validation of the proposed computational framework.
V Results
In this section, we present the results obtained when exploiting the EI and DI algorithms for detecting the EZ, considering the performance metrics given in Sec. IV-B. We remark that our computational framework is publicly available at the link: https://github.com/masonfed/desync_index. At the end of the section, we review in more detail two specific seizures among those analyzed, comparing the algorithm outcomes with clinical observations.
V-A Detection Results
Figure 1: Performance metrics in the EZ detection.
The SEEG signals are segmented in time windows lasting $T_{\text{window}}=1.00$ s each, considering a time shift of $\Delta t=0.25$ s. Given the epoch duration ($T_{\text{epoch}}=200$ s) a total of $797$ windows is analyzed for each seizure. For estimating the connectivity between different channels according to the PTE algorithm, we consider $\tau_{max}=0.10$ s as maximum propagation delay. Instead, for computing the energy ratio, the high- and low-frequency ranges are set to $B_{h}=[30.0,250.0]$ Hz and $B_{\ell}=[4.0,12.0]$ Hz, respectively. We observe that such a setting extends the portion of the spectrum processed by the EI algorithm, which, in its initial version, considered frequency lower than $97$ Hz because of the limited sampling rate [7]. In this work, we consider a wider frequency range to capture ripple phenomena that are notably characterized by oscillations up to $250$ Hz or even more [43]. On the other hand, this can lead to the inclusion of noisy information within our framework, leading to more false positive outcomes using either the EI and the DI algorithms.
In this work, we set the maximum number of epileptogenic channels to $M=15$, the tonicity interval to $\delta=5.00$ s (as done in [7]), the weight of the CUSUM chart to $\gamma=0.0$, and the exponential decay of the EWMA chart to $\alpha=0.1$. The tuning of $\gamma$ and $\alpha$ followed the recommendations provided in [44], and are kept to low values to capture the fast dynamics that characterized ictal phenomena. We set the detection threshold of the algorithms to $\eta=0.2$, which implies that we consider as part of the EZ only those channels presenting an epileptogenic value higher than $0.2$. Changing $\eta$ as well as $M$ affects the well-known trade-off between sensitivity and precision of detection algorithms. Finally, we set the starting time of the baseline period of the CUSUM chart to $t_{\text{base}}=t_{\text{start}}-50.0$ s. This implies that energy and connectivity statistics are estimating according to the signal evolution during the $50.0$ s preceding the seizure onset. As explained in Sec. IV, $t_{\text{start}}$ and $t_{\text{end}}$ depend on the specific seizure analyzed and, thus, are epoch-tailored.
(a) D-EZ.
(b) D-PZ.
Figure 2: Performance metrics in the D-EZ and D-PZ detection.
In Fig. 1, we report the detection performance of EI, DI, and the algorithm combinations, considering the channels within the EZ as a classification target. We can appreciate that DI shows higher sensitivity than EI ($0.39$ and $0.32$ in the two cases, respectively) when they are used as a standalone tool, while their combination (EI or DI) increases the sensitivity of more than $50$% (with a final score of $0.54$). This denotes that two algorithms return different sets of results, i.e., the channels associated with a higher EI value are not the same associated with a higher DI value. Focusing on precision, the EI algorithm leads to the best results (with a score of $0.51$), while considering the “and” function to combine the algorithms leads to a performance gain of $30\%$. On the other hand, the DI algorithm leads to a higher number of false positives and is indeed associated with a precision of only $0.35$. In general, the best accuracy is obtained using the “EI or DI” approach, which returns a score of $0.96$.
In Fig. 2, we report the detection performance considering the channels of the D-EZ and D-PZ zones as a classification target. We recall that D-EZ and D-PZ are defined according to the output of the “EI or DI” system, including all the epileptogenic and propagation channels that any of the algorithms classify as positive. In both cases, the DI algorithm returns the best results in terms of sensitivity ($0.73$ and $0.64$, respectively), while the EI algorithm leads to higher precision. This confirms that the use of DI, enables a more holistic analysis of the overall ictal transition, including more channels among those suspected of being part of the EZ, but may produce many false positives in the case of complex network topologies. On the other hand, DI is not intended to directly plan epilepsy surgery but to support neurophysiological evaluation. In a practical scenario, neurophysiologists themselves can discern false and true positives for the algorithm outcomes, making the sensitivity the primary performance indicator for our framework.
V-B Concordant Example
(a) Patient $7$.
(b) Patient $8$.
Figure 3: SEEG implants.
In the following, we better evaluate the operations of the two algorithms, analyzing two specific seizure events. We first consider patient $7$, whose SEEG implant (represented in Fig. 3(a)) explored the right temporo-parieto-occipital area, and includes $152$ different recording sites ($18$ electrodes). In Fig. 4(a), we report the bipolar representation of the $M=15$ channels associated with the strongest variation in the energy ratio during the 200 seconds preceding the propagation of epileptic discharges. Instead, in Fig. 4(b), we report the energy ratio values associated with each channel during the same period. In the two figures, the time is shifted according to the start of the ictal phase, so that $t_{\text{start}}=0$ s and $t_{\text{base}}=-50$ s. In particular, $t_{\text{base}}$ denotes the instant from which the baseline statistics of the energy ratio per channel were calculated, $t_{\text{base}}$ denotes the instant from which the algorithm starts detecting signal abnormalities, while the markers $t_{x}$ denote the activation times, i.e., the times at which the energy ratio $E_{x}$ diverges from its baseline.
(a) Signal.
(b) Energy ratio.
Figure 4: EI operation in a concordant case (patient $7$).
(a) Signal.
(b) Desynchronization level.
Figure 5: DI operation in a concordant case (patient $7$).
In Fig. 5, we represent the operation of the DI algorithm in the same scenario. Hence, Fig. 5(a) reports the $M=15$ channels showing the most significant connectivity changes, while Fig. 5(b)) reports the desynchronization level for each of the same channels. Comparing Fig. 5(b) with Fig. 4(b), we can observe that the energy ratio represents a less noisy metric than the desynchronization level. During the baseline, the energy ratio of each channel tends to zero, while, for $t_{\text{start}}>0$, most channels present abrupt energy changes, with a peak $E_{x}$ for channels Y4 and Y5. The desynchronization level is characterized by a much higher variability: many channels present bursts of reduced connectivity for $t_{\text{start}}<0$. On the other hand, Fig. 5(b) identifies desynchronization phenomena in cortical sites that, apparently, do not present variation in terms of fast oscillations. This is the case of channel V8, which starts desynchronizing from the rest of the network tens of seconds before the seizure onset but is not recognized as epileptogenic by the EI algorithm.
(a) EI.
(b) DI.
Figure 6: EI and DI outcomes in a concordant case (patient $7$).
Finally, in Fig. 6, we report the epileptogenicity levels assigned to each channel by the two algorithms, as well as the brain structure with which each channel is associated. In both cases, the output of the algorithms is normalized within the $[0,1]$ range, which means that epileptogenic values close to $1$ are associated with a greater influence in the diffusion of epileptic discharge. We can appreciate that the outputs are mostly concordant as both techniques identify the most epileptogenic channels within the posterior collateral sulcus and the fusiform gyrus. At the same time, DI is the sole tool that marks the channels within the the lingual gyrus, which was considered one of the trigger of the patient epilepsy.
V-C Discordant Example
(a) Signal.
(b) Energy ratio.
Figure 7: EI operation in a discordant case (patient $8$).
(a) Signal.
(b) Desynchronization level.
Figure 8: DI operation in a discordant case (patient $8$).
In the following, we consider patient $8$, whose SEEG implant (represented in Fig. 3(b)) explored the left fronto-temporo-parietal, area, and includes $116$ recording sites ($13$ electrodes). As done previously, we report the $M=15$ channels considered as most epileptogenic according to the two approaches, focusing first on the EI algorithm (Fig. 9(a)) and then on the DI algorithm (Fig. 9(b)). As before, the energy ratio associated with each channel presents low values during the baseline and then abruptly rises during the ictal phase. In particular, the channels associated with electrode I, crossing the temporal pole, maximize the energy ratio a few seconds after $t_{\text{start}}$.
In Fig. 9(b), we perform the same analysis for the desynchronization level. In this case, the network connectivity is less noisy than the seizure analyzed in Fig. 5 since the channels do not show prominent variations during the baseline period. This difference may be due to the different network sizes: the SEEG implant of patient $8$ includes $36$ more recording sites than that of patient $7$. In other words, in this case, the SEEG implant explores a much more focused epileptogenic network and, thus, leads to more stable connectivity dynamics.
Focus on the ictal phase ($t_{\text{start}}>0$), we can appreciate that DI identifies desynchronization phenomena in several structures that, instead, are not marked by epileptogenic in terms of energy ratio. As shown in Fig. 9(b), channels associated with electrodes B, C, and D (which all explore the cortical area neighboring the hyppocampus and parahyppocampus) are characterized by a significant drop in connectivity a few seconds after the start of the ictal phase. None of such channels (but B1) was included in those marked by EI, which can lead to the definition of a more restricted resection area during epilepsy surgery, with a lower chance of totally suppressing the mechanisms underlying seizure activation.
(a) EI.
(b) DI.
Figure 9: EI and DI outcomes in a discordant case (patient $8$).
Finally, Fig. 9 compares the epileptogenicity levels assigned to each channel by the algorithms, also reporting the cortical locations of the channels within the SEEG implants. In this case, the algorithm outputs is only partially concordant: EI marked the channels associated with the temporal pole as the most epileptogenic, while DI seems to identify the origin of the epileptic discharge in the hyppocampus. In particular, using EI, the most epileptogenic channel is I1, while the DI algorithm underscores channel C5, which is located within the hippocampal tail. This heterogeneous results denote how it is necessary to integrate multiple techniques, considering both energy-based and connectivity-based metrics, to improve the accuracy of EZ localization.
V-D Clinical Discussion
Our results show that the DI algorithm can recognize the desynchronization phenomena that precede the generation of the epileptic discharge in electrodes that intersect or surround the EZ. This leads DI to have a partially concordant output with EI, which, notably, aims at quantifying the magnitude of the high-frequency energy associated with ictal activities. Contrary to EI and most state-of-the-art solutions, DI does not compute the signal energy over specific frequency bands, but evaluates the desynchronization level of each SEEG channel considering all its oscillatory components. This makes it possible to implement the algorithm in a patient-agnostic fashion, without the need for assumptions about the frequency ranges characterizing ictal phenomena. Refining DI to consider specific frequency range is also possible and may further improve the accuracy of our methodology, although this may constitute a biasing factor for the results and reduce the generalizability of the framework.
When combined with EI, the proposed DI algorithm allows us to improve the accuracy in EZ definition and, used as a standalone tool, shows higher performance also in the detection of the cortical sites within the PZ. From a practical perspective, DI marks as epileptogenic the cortical sites that abruptly assume a static phase distribution in time and, thus, are not influenced by any other sites explored through the SEEG implant. Hence, the key idea behind our methodology is that epileptogenic signals present an independent behavior that cannot be inhibited by the other cortical structures. This is in line with the most recent findings in the field of EZ detection and confirms that the generation of epileptic seizures is allowed by a connectivity reduction in the overall network. Hence, during the inter-ictal phase, the regions surrounding the EZ may assume an inhibitory function that prevents the diffusion of epileptic discharges and, thus, constitute a key component of the epileptogenic network.
In a clinical context, the DI algorithm is not intended to directly plan the resection area for epilepsy surgery, but constitutes a support for neurophysiologists to identify the EZ and cortical sites that contribute to the propagation of epileptic discharges. In particular, DI emphasizes phenomena that are not marked by other computational techniques, such as EI, and are not easy to identify by a pure visual analysis. Our results prove that the combined use of DI and EI can strongly increase sensitivity in detecting ictal phenomena, although marking as epileptogenic cortical sites that are not part of the EZ. From this perspective, the clinical judgment of neurophysiologists remains an essential element to discern false and true positive, counterbalancing the lower precision of our framework. On the other hand, we remark that discerning ictal and non-ictal channels in a limited set, as the one given by our framework is a rather low-cost task, while reviewing all the SEEG channels is highly expensive and may lead to fatal human errors.
Finally, we observe that the proposed DI algorithm may also reflect the presence of ictal activities in cortical areas not directly explored by the SEEG implant. This may be useful in scenarios where the SEEG signal does not enable the clear visualization of the seizure generation and propagation. Hence, our technique may help neurophysiologists to discern whether the epileptogenic network is totally or only partially explored by the SEEG implant and, thus, the information obtained by the SEEG monitoring is sufficient to ensure, with high probability, the success of epilepsy surgery.
VI Conclusion
In this work, we developed a new computational framework for analyzing SEEG signals, with the final goal of identifying the EZ in patients with drug-resistant epilepsy. Our method, named DI, exploits the PTE model to estimate the effective connectivity between the cortical sites explored via SEEG and consider as most epileptogenic the SEEG channels that present an abrupt desynchronization in the period immediately preceding the seizure propagation. To evaluate its clinical utility, we implemented DI over a dataset of $11$ patients and compared its performance in identifying epileptogenic sites against the EI algorithm. Our results denoted that DI shows a higher sensitivity than EI in the detection of both the EZ and the PZ, and the best accuracy is obtained when combining the two algorithms as a unique detection tool.
The proposed approach, integrated with other quantitative biomarkers, may constitute a key support for SEEG interpretation. Indeed, DI underscores signal modifications not visually evident, helping to identify those sites that contribute to the seizure propagation or inhibit the epileptic discharge during the inter-ictal period. In future work, we intend to clinically validate the DI algorithm over a wider dataset, possibly including data from different clinical research centers, and evaluate the potential of our computational framework for analyzing SEEG signals during the inter-ictal period.